OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the second equation. 3x + y + 2z = 1 x + 2y – z = –8 3x – y – 3z = 18 –5y + 5z = 25 –7y = 42 –5x – 5z = –10 6x – z = 19 7x + 3z = –10 0x – 5z = 19 7y – z = 25 5y – 6z = 42 If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the second equation. 3x + y + 2z = 1 x + 2y – z = –8 3x – y – 3z = 18 –5y + 5z = 25 –7y = 42 –5x – 5z = –10 6x – z = 19 7x + 3z = –10 0x – 5z = 19 7y – z = 25 5y – 6z = 42 @Mathematics

OpenStudy (radar):

Step 1. Isolate x in second equation. x=-8+z-2y using algebra. Step 2. substitute this new value of x in first and third equation. First equation becomes 3(-8+z-2y)+y+2z=1 Third equation becomes 3(-8+z-2y)-y-3z=18 Step 3. Now simplify those equations. first one first:-24+3z-6y+y+2z=1 -5y+5z=25 third one next -24+3z-6y-y-3z=18 -7y=42 There you have it: -5y+5z=25 and -7y=42 The first two.

OpenStudy (anonymous):

radar is correct.

OpenStudy (radar):

Hello robtobey, how are things in CA?

OpenStudy (anonymous):

Raining today. Otherwise everything is normal.

OpenStudy (radar):

It has turned of windy and chilly, but lots of sunshine.

OpenStudy (anonymous):

Where are you located again?

OpenStudy (radar):

the Ozarks of MO

OpenStudy (anonymous):

What is the highest elevation in your immediate area?

OpenStudy (radar):

Tom Sauk Mt. however only 1772 ft. Doesn't compare to the sierrias or cascades etc.

OpenStudy (anonymous):

Just checked out Truckee, CA at 5829 ft and it is snowing there today.

OpenStudy (radar):

Is that up near Lake Tahoe?

OpenStudy (anonymous):

It is on interstate 80 on the way to Reno, NV

OpenStudy (radar):

going to pop some popcorn and watch some TV with my wife. CUL

OpenStudy (anonymous):

Nice talking to you.

OpenStudy (radar):

checking out you link. Good to hear from you.